# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v4

@@ Appendix A Theoretical Foundations of Task Entropy
This section provides the theoretical foundations of task entropy, quantifying the complexity of learning tasks. The concept of task entropy was first proposed for the image level and formulated as Li (2022):
$H(\mathcal{T};\bm{X})=-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X})\log{p(\bm{Y}|\bm{X})}$ (22)
The image $\bm{X}$ in the dataset are supposed to be independent of each other, as are the labels $\bm{Y}$. However, $\bm{X}$ and $\bm{Y}$ are not independent because of the correlation between a data sample $X$ and its corresponding label $Y$. Essentially, the task entropy is the entropy of $p(Y|X)$. Following the principle of task entropy, compelling evidence suggests that diminishing task complexity via reducing information entropy can enhance overall model performance Li (2022); Zhang et al. (2023).
@@
$H(\mathcal{T};\bm{Z})\coloneqq H(\bm{Y},\bm{Z})-H(\bm{Z})$ (23)
where $\bm{Z}$ are the embeddings of the images $\bm{X}$. Here, we assume that the embedding $\bm{Z}$ and the vectorized label $\bm{Y}$ follow a multivariate normal distribution. We can transform the unknown distributions of $\bm{Z}$ and $\bm{Y}$ to approximately conform to normality by utilizing techniques such as Box-Cox transformation, log transform, etc., as described in Box & Cox (1964) and Feng et al. (2014) if they are not initially normal. After approximate transformation, the distribution of $\bm{Z}$ and $\bm{Y}$ can be expressed as:
$\bm{Z}\sim\mathcal{N}(\bm{\mu_{Z}},\bm{\Sigma_{Z}}),\bm{Y}\sim\mathcal{N}(\bm{\mu_{Y}},\bm{\Sigma_{Y}})$ (24)
@@
$\begin{split}&\bm{{\mu}_{Z}}=\mathbb{E}[\bm{Z}]=(\mathbb{E}[Z_{1}],\mathbb{E}[Z_{2}],...,\mathbb{E}[Z_{k}]])^{T}\\ &\bm{{\mu}_{Y}}=\mathbb{E}[\bm{Y}]=(\mathbb{E}[Y_{1}],\mathbb{E}[Y_{2}],...,\mathbb{E}[Y_{k}]])^{T}\\ &\bm{\Sigma_{Z}}=\mathbb{E}[(\bm{Z}-\bm{{\mu}_{Z}})(\bm{Z}-\bm{{\mu}_{Z}})^{T}]\\ &\bm{\Sigma_{Y}}=\mathbb{E}[(\bm{Y}-\bm{{\mu}_{Y}})(\bm{Y}-\bm{{\mu}_{Y}})^{T}]\end{split}$ (25)
@@
Then the conditional distribution of $\bm{Y}$ given $\bm{Z}$ is also normally distributed Mood (1950) Johnson et al. (1995), which can be formulated as:
$\bm{Y}|\bm{Z}\sim\mathcal{N}(\mathbb{E}({\bm{Y}|\bm{Z}={Z}}),var(\bm{Y}|\bm{Z}={Z}))$ (26)
where $\mathbb{E}({\bm{Y}|\bm{Z}=Z})$ is the mean of the label set $\bm{Y}$ given a sample $\bm{Z}=Z$ from the embeddings, and $var(\bm{Y}|\bm{Z}=Z)$ is the variance of $\bm{Y}$ given a sample from the embeddings. The conditional mean $\mathbb{E}[(\bm{Y}|\bm{Z}=Z)]$ and conditional variance $var(\bm{Y}|\bm{Z}=Z)$ can be calculated as:
$\bm{\mu}_{\bm{Y}|\bm{Z}=Z}=\mathbb{E}[(\bm{Y}|\bm{Z}=Z)]=\bm{{\mu}_{Y}}+\bm{\Sigma_{YZ}}\bm{{\Sigma}^{-1}_{Z}}(Z-\bm{\mu_{Z}})$ (27)
$\bm{\Sigma}_{{\bm{Y|Z}}=Z}=var(\bm{Y}|\bm{Z}=Z)=\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}$ (28)
where $\bm{\Sigma_{YZ}}$ and $\bm{\Sigma_{ZY}}$ are the cross-covariance matrices between $\bm{Y}$ and $\bm{Z}$, and between $\bm{Z}$ and $\bm{Y}$, respectively, and $\bm{\Sigma^{-1}_{Z}}$ denotes the inverse of the covariance matrix of $\bm{Z}$.
@@
$\begin{split}H(\mathcal{T};\bm{Z})=&-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{Z})\log{p(\bm{Y}|\bm{Z})}\\ =&-\mathbb{E}[\log p(\bm{Y}|\bm{Z})]\\ =&-\mathbb{E}\bigg{[}\log\big{[}(2\pi)^{-k/2}|\bm{\Sigma_{Z}}|^{-1/2}\\ &\quad\times\exp\big{(}-\frac{1}{2}(\bm{Y|Z}-\bm{\mu_{Y|Z}})^{T}\bm{\Sigma_{Y|Z}}^{-1}(\bm{Y|Z}-\bm{\mu_{Y|Z}})\big{)}\big{]}\bigg{]}\\ =&\frac{k}{2}\log(2\pi)+\frac{1}{2}\log|\bm{\Sigma_{Y|Z}}|\\ &+\frac{1}{2}\mathbb{E}[(\bm{Y|Z}-\bm{\mu}_{\bm{Y|Z}})^{T}\bm{\Sigma_{Y|Z}}^{-1}(\bm{Y|Z}-\bm{\mu}_{\bm{Y|Z}})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\bm{\Sigma_{Y|Z}}|\end{split}$ (29)
@@
$\begin{split}\mathbb{E}\left[(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\Sigma^{-1}_{\bm{Y\mid Z}}(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})\right]=&\mathbb{E}\left[\mathrm{tr}\left((\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\Sigma^{-1}_{\bm{Y\mid Z}}(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})\right)\right]\\ =&\mathbb{E}\left[\mathrm{tr}\left(\Sigma^{-1}_{\bm{Y\mid Z}}(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\right)\right]\\ =&\mathrm{tr}\left(\Sigma^{-1}_{\bm{Y\mid Z}}\mathbb{E}\left[(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\right]\right)\\ =&\mathrm{tr}\left(\Sigma^{-1}_{\bm{Y\mid Z}}\Sigma_{\bm{Y\mid Z}}\right)\\ =&\mathrm{tr}(\bm{I}_{k})\\ =&k\end{split}$ (30)
Therefore, for a specific set of embeddings, we can find that the task entropy is only related to the variance of the $\bm{Y|Z}$.
@@
$\left\{\begin{matrix}H(\mathcal{T};\bm{Z}+\bm{\epsilon})\coloneqq H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})&\bm{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\bm{Z}\bm{\epsilon})\coloneqq H(\bm{Y};\bm{Z}\bm{\epsilon})-H(\bm{Z})&\bm{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (31)
@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{matrix}H(\mathcal{T};\bm{Z}+\bm{\epsilon})=H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})\end{matrix}.$ (32)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&\ H(\mathcal{T};\bm{Z})-H(\mathcal{T};\bm{Z}+\bm{\epsilon})\\ =&\ H(\bm{Y};\bm{Z})-H(\bm{Z})-\left(H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})\right)\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}+\bm{\epsilon})\\ =&H(\bm{Y}|\bm{Z})+H(\bm{Z})-\left(H(\bm{Y}|\bm{Z}+\bm{\epsilon})+H(\bm{Z}+\bm{\epsilon})\right)\\ =&\frac{1}{2}\log|\bm{\Sigma_{Y|Z}}|+\frac{1}{2}\log|\bm{\Sigma_{Z}}|\\ &-\frac{1}{2}\log|\bm{\Sigma_{Y|Z+\epsilon}}|-\frac{1}{2}\log|\bm{\Sigma_{Z+\bm{\epsilon}}}|\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y}}-\Sigma_{YZ}\Sigma_{Z}^{-1}\Sigma_{ZY}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y}}-\Sigma_{YZ}\Sigma_{Z+\epsilon}^{-1}\Sigma_{ZY}|}\end{split}$ (33)
where $\bm{\Sigma_{Y|Z+\epsilon}}=\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}$. Since the Gaussian noise is independent of $\bm{Z}$ and $\bm{Y}$, we have $\bm{\Sigma_{Y(Z+\epsilon)}}=\bm{\Sigma_{(Z+\epsilon)Y}}=\bm{\Sigma_{YZ}}$. The corresponding proof is:
$\begin{split}\bm{\Sigma_{(Z+\epsilon)Y}}=&\mathbb{E}\left[\bm{(Z+\epsilon)}-\mu_{\bm{Z+\epsilon}}\right]\mathbb{E}\left[\bm{Y}-\mu_{\bm{Y}}\right]\\ =&\mathbb{E}\left[\bm{(Z+\epsilon)}\bm{Y}\right]-\mu_{\bm{Y}}\mathbb{E}\left[\bm{(Z+\epsilon)}\right]-\mu_{\bm{Z+\epsilon}}\mathbb{E}\left[\bm{Y}\right]+\mu_{\bm{Y}}\mu_{\bm{Z+\epsilon}}\\ =&\mathbb{E}\left[\bm{(Z+\epsilon)}\bm{Y}\right]-\mu_{\bm{Y}}\mathbb{E}\left[\bm{(Z+\epsilon)}\right]\\ =&\mathbb{E}\left[\bm{Z}\bm{Y}\right]+\mathbb{E}\left[\bm{\epsilon}\bm{Y}\right]-\mu_{\bm{Y}}\mu_{\bm{Z}}-\mu_{\bm{Y}}\mu_{\bm{\epsilon}}\\ =&\mathbb{E}\left[\bm{Z}\bm{Y}\right]-\mu_{\bm{Y}}\mu_{\bm{Z}}\\ =&\bm{\Sigma_{ZY}}\end{split}$ (34)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\bm{\epsilon})>0&if\ \frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|}>1\\ \bigtriangleup S(\mathcal{T},\bm{\epsilon})\leq 0&if\ \frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|}\leq 1\end{matrix}\right.$ (35)
To find the relationship between $|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|$ and $|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|$, we need to determine the subterms in each of them. As we mentioned in the previous section, the embeddings of the images are independent of each other, and so are the labels.
$\begin{split}\bm{\Sigma_{Y}}=&\mathbb{E}[(\bm{Y}-\bm{{\mu}_{Y}})(\bm{Y}-\bm{{\mu}_{Y}})^{T}]\\ =&\mathbb{E}[\bm{Y}\bm{Y}^{T}]-\bm{{\mu}_{Y}}\bm{{\mu}_{Y}}^{T}\\ =&\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})\end{split}$ (36)
@@
The same procedure can be applied to $\bm{\Sigma_{Y(Z+\epsilon)}}$ and $\bm{{\Sigma}_{Z+\epsilon}}$. Therefore, We can obtain that $\bm{\Sigma_{Y}}=\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})$,
$\bm{\Sigma_{Y(Z+\epsilon)}}=\mathrm{diag}(\mathrm{cov}(Y_{1},Z_{1}+\epsilon),...,\mathrm{cov}(Y_{k},Z_{k}+\epsilon))$ (38)
and $\bm{{\Sigma}_{Z+\epsilon}}$ is:
@@
$\begin{split}\bm{{\Sigma}_{Z+\epsilon}}=&\begin{bmatrix}\sigma^{2}_{Z_{1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{2}}+\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \vdots&\vdots&&\vdots&\vdots\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{Z_{k-1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{k}}+\sigma^{2}_{\epsilon}\\ \end{bmatrix}\\ =&\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\bm{I}_{k}+\sigma^{2}_{\epsilon}\bm{1}_{k}\end{split}$ (39)
@@
where $\bm{I}_{k}$ is a $k\times k$ identity matrix and $\bm{1}_{k}$ is a all ones $k\times k$ matrix. We use $\bm{U}$ to represent $\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\bm{I}_{k}$, and $\bm{u}$ to represent a all ones vector $[1,...,1]^{T}$. Thanks to the Sherman–Morrison Formula Sherman & Morrison (1949) and Woodbury Formula Woodbury (1950), we can obtain the inverse of $\bm{{\Sigma}_{Z+\epsilon}}$ as:

### v5

@@ Appendix A Theoretical Foundations of Task Entropy
This section provides the theoretical foundations of task entropy, quantifying the complexity of learning tasks. The concept of task entropy was first proposed for the image level and formulated as [30]:
$H_{\mathcal{T}}(\boldsymbol{X})=-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X})\log{p(\boldsymbol{Y}|\boldsymbol{X})}$ (21)
The image $\boldsymbol{X}$ in the dataset are supposed to be independent of each other, as are the labels $\boldsymbol{Y}$. However, $\boldsymbol{X}$ and $\boldsymbol{Y}$ are not independent because of the correlation between a data sample $X$ and its corresponding label $Y$. Essentially, the task entropy is the entropy of $p(Y|X)$. Following the principle of task entropy, compelling evidence suggests that diminishing task complexity via reducing information entropy can enhance overall model performance [30, 69].
@@
$H_{\mathcal{T}}(\boldsymbol{Z})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})$ (22)
where $\boldsymbol{Z}$ are the embeddings of the images $\boldsymbol{X}$. Here, we assume that the embedding $\boldsymbol{Z}$ and the vectorized label $\boldsymbol{Y}$ follow a multivariate normal distribution. We can transform the unknown distributions of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ to approximately conform to normality by utilizing techniques if they are not initially normal. After approximate transformation, the distribution of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ can be expressed as:
$\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\mu_{Z}},\boldsymbol{\Sigma_{Z}}),\boldsymbol{Y}\sim\mathcal{N}(\boldsymbol{\mu_{Y}},\boldsymbol{\Sigma_{Y}})$ (23)
@@
$\begin{split}&\boldsymbol{{\mu}_{Z}}=\mathbb{E}[\boldsymbol{Z}]=(\mathbb{E}[Z_{1}],\mathbb{E}[Z_{2}],...,\mathbb{E}[Z_{k}]])^{T}\\ &\boldsymbol{{\mu}_{Y}}=\mathbb{E}[\boldsymbol{Y}]=(\mathbb{E}[Y_{1}],\mathbb{E}[Y_{2}],...,\mathbb{E}[Y_{k}]])^{T}\\ &\boldsymbol{\Sigma_{Z}}=\mathbb{E}[(\boldsymbol{Z}-\boldsymbol{{\mu}_{Z}})(\boldsymbol{Z}-\boldsymbol{{\mu}_{Z}})^{T}]\\ &\boldsymbol{\Sigma_{Y}}=\mathbb{E}[(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})^{T}]\end{split}$ (24)
@@
Then the conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{Z}$ is also normally distributed [38] [24], which can be formulated as:
$\boldsymbol{Y}|\boldsymbol{Z}\sim\mathcal{N}(\mathbb{E}({\boldsymbol{Y}|\boldsymbol{Z}={Z}}),var(\boldsymbol{Y}|\boldsymbol{Z}={Z}))$ (25)
where $\mathbb{E}({\boldsymbol{Y}|\boldsymbol{Z}=Z})$ is the mean of the label set $\boldsymbol{Y}$ given a sample $\boldsymbol{Z}=Z$ from the embeddings, and $var(\boldsymbol{Y}|\boldsymbol{Z}=Z)$ is the variance of $\boldsymbol{Y}$ given a sample from the embeddings. The conditional mean $\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{Z}=Z)]$ and conditional variance $var(\boldsymbol{Y}|\boldsymbol{Z}=Z)$ can be calculated as:
$\boldsymbol{\mu}_{\boldsymbol{Y}|\boldsymbol{Z}=Z}=\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{Z}=Z)]=\boldsymbol{{\mu}_{Y}}+\boldsymbol{\Sigma_{YZ}}\boldsymbol{{\Sigma}^{-1}_{Z}}(Z-\boldsymbol{\mu_{Z}})$ (26)
$\boldsymbol{\Sigma}_{{\boldsymbol{Y|Z}}=Z}=var(\boldsymbol{Y}|\boldsymbol{Z}=Z)=\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}$ (27)
where $\boldsymbol{\Sigma_{YZ}}$ and $\boldsymbol{\Sigma_{ZY}}$ are the cross-covariance matrices between $\boldsymbol{Y}$ and $\boldsymbol{Z}$, and between $\boldsymbol{Z}$ and $\boldsymbol{Y}$, respectively, and $\boldsymbol{\Sigma^{-1}_{Z}}$ denotes the inverse of the covariance matrix of $\boldsymbol{Z}$.
@@
$\begin{split}H_{\mathcal{T}}(\boldsymbol{Z})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{Z})\log{p(\boldsymbol{Y}|\boldsymbol{Z})}\\ =&-\mathbb{E}[\log p(\boldsymbol{Y}|\boldsymbol{Z})]\\ =&-\mathbb{E}[\log[(2\pi)^{-k/2}|\boldsymbol{\Sigma_{Z}}|^{-1/2}\exp(-\frac{1}{2}\boldsymbol{(Y|Z-\mu_{Y|Z})}^{T}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{Y|Z}}(\boldsymbol{Y|Z}-\boldsymbol{\mu_{Y|Z}}))]]\\ =&\frac{k}{2}\log(2\pi)+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|+\frac{1}{2}\mathbb{E}[{(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|\end{split}$ (28)
@@
$\begin{split}\mathbb{E}[{(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})]=&\mathbb{E}[tr((\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}))]\\ =&\mathbb{E}[tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}){(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T})]\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}){(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T})\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}\Sigma_{\boldsymbol{\boldsymbol{Y|Z}}})\\ =&tr(\boldsymbol{I}_{k})\\ =&k\end{split}$ (29)
Therefore, for a specific set of embeddings, we can find that the task entropy is only related to the variance of the $\boldsymbol{Y|Z}$.
@@
$\left\{\begin{matrix}H_{\mathcal{T}}(\boldsymbol{Z}+\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H_{\mathcal{T}}(\boldsymbol{Z}\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right$. (30)
@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{matrix}H_{\mathcal{T}}(\boldsymbol{Z}+\boldsymbol{\epsilon})=H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})\end{matrix}.$ (31)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H_{\mathcal{T}}(\boldsymbol{Z})-H_{\mathcal{T}}(\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z}))\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|+\frac{1}{2}\log|\boldsymbol{\Sigma_{Z}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Z+\boldsymbol{\epsilon}}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}\end{split}$ (32)
where $\boldsymbol{\Sigma_{Y|Z+\epsilon}}=\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}$. Since the Gaussian noise is independent of $\boldsymbol{Z}$ and $\boldsymbol{Y}$, we have $\boldsymbol{\Sigma_{Y(Z+\epsilon)}}=\boldsymbol{\Sigma_{(Z+\epsilon)Y}}=\boldsymbol{\Sigma_{YZ}}$. The corresponding proof is:
$\begin{split}\boldsymbol{\Sigma_{(Z+\epsilon)Y}}=&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}-\mu_{\boldsymbol{Z+\epsilon}}\right]\mathbb{E}\left[\boldsymbol{Y}-\mu_{\boldsymbol{Y}}\right]\\ =&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\right]-\mu_{\boldsymbol{Z+\epsilon}}\mathbb{E}\left[\boldsymbol{Y}\right]+\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z+\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\right]\\ =&\mathbb{E}\left[\boldsymbol{Z}\boldsymbol{Y}\right]+\mathbb{E}\left[\boldsymbol{\epsilon}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z}}-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{Z}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z}}\\ =&\boldsymbol{\Sigma_{ZY}}\end{split}$ (33)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&if\ \frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}>1\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&if\ \frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}\leq 1\end{matrix}\right$. (34)
To find the relationship between $|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|$ and $|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|$, we need to determine the subterms in each of them. As we mentioned in the previous section, the embeddings of the images are independent of each other, and so are the labels.
$\begin{split}\boldsymbol{\Sigma_{Y}}=&\mathbb{E}[(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})^{T}]\\ =&\mathbb{E}[\boldsymbol{Y}\boldsymbol{Y}^{T}]-\boldsymbol{{\mu}_{Y}}\boldsymbol{{\mu}_{Y}}^{T}\\ =&\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})\end{split}$ (35)
@@
The same procedure can be applied to $\boldsymbol{\Sigma_{Y(Z+\epsilon)}}$ and $\boldsymbol{{\Sigma}_{Z+\epsilon}}$. Therefore, We can obtain that $\boldsymbol{\Sigma_{Y}}=\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})$,
$\boldsymbol{\Sigma_{Y(Z+\epsilon)}}=\mathrm{diag}(\mathrm{cov}(Y_{1},Z_{1}+\epsilon),...,\mathrm{cov}(Y_{k},Z_{k}+\epsilon))$ (37)
and $\boldsymbol{{\Sigma}_{Z+\epsilon}}$ is:
@@
$\begin{split}\boldsymbol{{\Sigma}_{Z+\epsilon}}=&\begin{bmatrix}\sigma^{2}_{Z_{1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{2}}+\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \vdots&\vdots&&\vdots&\vdots\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{Z_{k-1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{k}}+\sigma^{2}_{\epsilon}\\ \end{bmatrix}\\ =&\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\boldsymbol{I}_{k}+\sigma^{2}_{\epsilon}\boldsymbol{1}_{k}\end{split}$ (38)
@@
where $\boldsymbol{I}_{k}$ is a $k\times k$ identity matrix and $\boldsymbol{1}_{k}$ is a all ones $k\times k$ matrix. We use $\boldsymbol{U}$ to represent $\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\boldsymbol{I}_{k}$, and $\boldsymbol{u}$ to represent a all ones vector $1,...,1]^{T}$. Thanks to the Sherman–Morrison Formula [51] and Woodbury Formula [64], we can obtain the inverse of $\boldsymbol{{\Sigma}_{Z+\epsilon}}$ as: